OpenStudy (anonymous):

Which of the following is a zero for the function f(x) = (x + 3)(x - 7)(x + 5)?

OpenStudy (anonymous):

arent there choices? easiest way since these are all multiplication is to set each expression equal to 0 and solve for x.

OpenStudy (anonymous):

Select one: A. x = -7 B. x = -3 C. x = 3 D. x = 5 Oops! Sorry! I forgot to put them down! Thank you!

OpenStudy (anonymous):

ok. I can't just give you the answer, but like I said, since these are all multiplication, just take each term and set it equal to zero.. example (x+3)=0 .. now solve for x. so this function will have 3 zeros.

OpenStudy (anonymous):

because if you multiply anything by 0 its 0 right?

OpenStudy (anonymous):

Ok, so you're saying that all of them can be the answer? I'm confused?

OpenStudy (anonymous):

what i am saying is that this function has 3 terms.. do we agree?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok. now, if you set each of those separately = to 0 and solve for x you can get your zeros. (x + 3)(x - 7)(x + 5) so lets to x-7 x-7=0.. add 7 to both sides. x = 7

OpenStudy (anonymous):

so 7 is a zero of this function

OpenStudy (anonymous):

(x + 3) x=3 (x - 7) x=7 (x + 5) x=5

OpenStudy (anonymous):

3+3 is not 0 and either is 5+5

OpenStudy (anonymous):

(x + 3) x=6 (x + 5) x=10

OpenStudy (anonymous):

x+5=0 so...what plus 5 equals 0?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

exactly. so -5 is also a zero of the function.

OpenStudy (anonymous):

now just one term left!

OpenStudy (anonymous):

x+3=0 -3+3=0

OpenStudy (anonymous):

you got it man. awesome! do you understand the concept? or can I be of more help?

OpenStudy (anonymous):

Sorry for guessing! I'm just confused on this subject! Thank you so much for your help!

OpenStudy (anonymous):

its my pleasure to help. hope you understand, because thats all the teacher really wants.

OpenStudy (anonymous):

Ok, so I get the concept, it's just I don't understand how to answer this question.

OpenStudy (anonymous):

well what are the zeros? -5, 7, and -3? so if anyone of these 3 terms is equal to zero.. meaning then the function is equal to zero.

OpenStudy (anonymous):

only because we are doing (term1) x (term2) x (term3). so if x=5, then term 3 = 0 and no matter what the other terms equal.. (term1)(term2)(zero) the function is still equal to zero.. only because we are multiplying.

OpenStudy (anonymous):

the zeros simply means what number can x be where the function will be equal to zero.

OpenStudy (anonymous):

and you already determined the "Zeros" of the function. x=-5, x=7, and x=-3

OpenStudy (anonymous):

Alright! Cool! So I'm guessing that the answer is B. x= -3. You don't have to tell me if that answer is correct if it is against the rules of this website. I've just joined openstudy today, obviously, so I'm clueless as to what the rules and regulations are.

OpenStudy (anonymous):

nah, i just dont give straight answers. and you dont have to guess, because you already figured out that x = -3 is a zero of the function f(x)= (x + 3)(x - 7)(x + 5) this is true because (-3+3=0)(-3-7=-10)(-3+5=2) =0

OpenStudy (anonymous):

i just want to help people learn, and i figure it doesnt do you any good for me to just give you the answer.. especially with math.. since you need todays work for tomorrow's assignment.

OpenStudy (anonymous):

let me know if you need any more help, I will be around for a few hours. working on homework myself.

OpenStudy (anonymous):

I 100% agree with you! Thank you very much for helping me understand this! I will probably be around with a few more questions! Thanks again! :)